Can you please explain how to solve for S in this problem, it's been a very long time since I've had to do problems of this nature. Thanks
2[s+3(s-1)]=18

Answer by rfer(16322)

work your way out of thr parentheses and brackets first
2[s+3(s-1)]=18
2[s+3s-3]=18
2[4s-3]=18
8s-6=18
add 6 to both sides
8s=24
s=3